OpenStudy (anonymous):

I need someone to explain how to do this.. How do you write an equation for the nth term of the Arithmetic Sequence 4, 1, -2, -5? I know the difference is -3 and a1 is 4 (i think) but what do i use that for and how do i write it and what is it?

OpenStudy (whpalmer4):

the formula to use is \[a_n = a_1 + (n-1)d\] where \(n\) is the number of the term you seek, \(d\) is the common difference, and \(a_1\) is the value of the first term.

OpenStudy (anonymous):

so would the equation just be An= 4 + (n-1) (-3)

OpenStudy (whpalmer4):

Let's find out! \[a_2 = a_1 + (2-1)(-3) = 4 + 1(-3) = 4-3 = 1\]\[a_3 = a_1+(3-1)(-3) = 4 +2(-3) = 4-6 = -2\]\[a_4 = 4+3(-3) = 4-9 = -5\] Does that look like it produces the right results?

OpenStudy (anonymous):

Oh! so i don't need anything else? i thought there was more to it!

OpenStudy (anonymous):

it came out with the right answers when i did it too!

OpenStudy (whpalmer4):

That's all there is to that problem, but there are some similar problems that we could do so you know how to do them as well. For example, what if you know the 1st and 3rd terms only? Could you figure out the formula?

OpenStudy (anonymous):

i dont know how you'd do that..

OpenStudy (whpalmer4):

Well, let's try it with a new sequence. 1, ???, 5, ???, etc. First, what's the value of \(a_1\)?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Good. So, what's the formula going to be (just use letters where you don't know the values)?

OpenStudy (anonymous):

\[An _{?} = 1 (n-1) (d)\]

OpenStudy (anonymous):

1+ (n-1) (d)*

OpenStudy (whpalmer4):

very good! by the way, to do a subscript in the equation editor, you write _ before the thing you want as a subscript. so a_n = a_1 + (n-1)d is what I type in. If you need a subscript to have more than one digit, you wrap it in { }, so \(a_{23}\) is produced with a_{23}

OpenStudy (whpalmer4):

So we have \[a_n = 1+(n-1)d\]Now the third term of the sequence is 5. Can we use that information to figure out the value of \(d\)?

OpenStudy (anonymous):

yes but i dont know how..

OpenStudy (whpalmer4):

Well, what is the formal name for the 3rd term?

OpenStudy (anonymous):

if theres an equation ment to find just the 3rd term my teacher didnt show it to us..

OpenStudy (whpalmer4):

oh, you know it, you figured it out in the other problem :-) \(a_3\) is the third term, isn't it? The sequence goes \(a_1, a_2, a_3, a_4, ..., a_n\)

OpenStudy (whpalmer4):

So we can write \[a_3 = a_1 + (n-1)d\]but if we are looking at \(a_n = a_3\) then we have \(n = 3\), right? And we already know that \(a_1 = 1\) so that gives us\[a_3 = 1 + (3-1)d\] Now can you find the value of \(d\)?

OpenStudy (anonymous):

oh oh! yeah i knew that.. and so a_3 = 1+ (4) d

OpenStudy (whpalmer4):

no, not quite...

OpenStudy (anonymous):

wait im very confused

OpenStudy (whpalmer4):

\[a_3 = 1 +(3-1)d\]\[a_3 = 1 + 2d\]\[5 = 1 + 2d\]right?

OpenStudy (whpalmer4):

I think you added instead of subtracting...

OpenStudy (whpalmer4):

you were excited :-)

OpenStudy (anonymous):

oh im sorry im not thinking straight and i did..

OpenStudy (whpalmer4):

Okay, so \(d = \)

OpenStudy (anonymous):

d is 2?

OpenStudy (anonymous):

so the equation would be A1= 1 + (n-1) (2)?

OpenStudy (whpalmer4):

Yes! So our sequence is \[a_n = 1 + (n-1)2\]or more usually \[a_n = 1+2(n-1)\](writing it the other way on OpenStudy will make people think that maybe you copied and pasted from something like \((n-1)^2\) and ended up with \((n-1)2\) which wasn't true)

OpenStudy (whpalmer4):

Okay, now for the final variation :-) What if I told you that you needed to find the formula for this sequence? ???, 4, ???, ???, 13

OpenStudy (anonymous):

2nd and 5th place.. hold on ill do it

OpenStudy (whpalmer4):

I've got a nice, shiny medal ready to award to you when you get it :-)

OpenStudy (anonymous):

do i need to find all of the empty spaces or just one?

OpenStudy (whpalmer4):

find the formula, and I'll take it for granted that you could find any space I asked for...

OpenStudy (anonymous):

i am stuck already..

OpenStudy (whpalmer4):

well, write out what you can, and I'll help you along.

OpenStudy (anonymous):

so you start with An= a1 + (n-1) (d)

OpenStudy (whpalmer4):

yes.

OpenStudy (whpalmer4):

but this time, you don't know the value of \(a_1\), do you?

OpenStudy (anonymous):

no..

OpenStudy (whpalmer4):

doesn't matter :-)

OpenStudy (whpalmer4):

write out what you do know for the two terms you know.

OpenStudy (anonymous):

a2= a1 + (2-1) (d)

OpenStudy (whpalmer4):

it's going to turn out to be essentially the same problem as finding the y-intercept of a line given two points on the line

OpenStudy (anonymous):

wait is that part right?

OpenStudy (whpalmer4):

yes, so far so good...

OpenStudy (anonymous):

now what?

OpenStudy (whpalmer4):

do the same with the other term you know...

OpenStudy (anonymous):

ok i did

OpenStudy (whpalmer4):

so what are the two equations you have, with all the numbers you know substituted in?

OpenStudy (anonymous):

the one i said already and a5= a1 + (5-1)(d)

OpenStudy (whpalmer4):

okay, but you haven't substituted in ALL of the numbers you know...

OpenStudy (anonymous):

what others do i know??

OpenStudy (whpalmer4):

don't you know the actual values of \(a_2\) and \(a_5\)?

OpenStudy (anonymous):

oh yeah do i plug that in where a1 is?

OpenStudy (whpalmer4):

No, you plug it in where its name is... if \(a_2 = 4\), then the equation is \[4 = a_1+(2-1)d\]\[4=a_1 + d\] Do the same for \(a_5\), please...

OpenStudy (anonymous):

13= a1+ (4) (d)

OpenStudy (whpalmer4):

So we have two equations: \[4 = a_1 +d\]\[13 = a_1 + 4d\] Is it possible to find the values of \(a_1\) and \(d\) from that alone?

OpenStudy (anonymous):

i mean i bet it is but i dont know how..

OpenStudy (whpalmer4):

Haven't you done systems of equations yet? Substitution, or elimination?

OpenStudy (anonymous):

not really no..

OpenStudy (whpalmer4):

Hmm, I thought that usually happened before arithmetic sequences, but maybe not. Okay, well, can you solve \[4 = a_1 +d\]for \(a_1\)? You won't get a number, you'll get an equation, but with only \(a_1\) on one side of the = sign.

OpenStudy (whpalmer4):

\[4 = a_1 + d\]Subtract \(d\) from both sides:\[4-d = a_1 + d - d\]collect like terms:\[4-d = a_1 \cancel{+d}\cancel{-d}\]\[a_1 = 4-d\]right?

OpenStudy (anonymous):

yes..

OpenStudy (whpalmer4):

Okay. Now take the second equation, and replace \(a_1\) with \(4-d\). What do you get?

OpenStudy (anonymous):

i was doing it wrong i got it now hold on!

OpenStudy (anonymous):

i am messing it up i think..

OpenStudy (anonymous):

which do i do first? subtract the 4 or divide the other 4?

OpenStudy (whpalmer4):

\[13 = a_1+4d\]and we found that \[a_1 = 4-d\] \[13 = (4-d) + 4d\] Can you solve that for \(d\)?

OpenStudy (anonymous):

divide the four first right

OpenStudy (whpalmer4):

No, get rid of the parentheses first.

OpenStudy (whpalmer4):

that gives us \[13 = 4-d+4d\] what is \[-d + 4d\]? Hint: \[-d = -1d\]

OpenStudy (anonymous):

9= -d+4d

OpenStudy (anonymous):

i subtracted the 4 from 13

OpenStudy (whpalmer4):

Good so far, keep going

OpenStudy (anonymous):

do i divide by 4?

OpenStudy (whpalmer4):

No. what does \(-d + 4d=\)

OpenStudy (anonymous):

9

OpenStudy (whpalmer4):

no, just simplify that expression...

OpenStudy (anonymous):

oh... wait you can't i didnt think

OpenStudy (whpalmer4):

sure you can. it's just the same as \[4d - d\] or \[4d - 1d\]

OpenStudy (anonymous):

oh!

OpenStudy (whpalmer4):

is there a big dent in your forehead now? :-)

OpenStudy (anonymous):

yeahh..

OpenStudy (whpalmer4):

so \[9 = -d+4d\]\[d = \]

OpenStudy (anonymous):

i am still confused..

OpenStudy (anonymous):

so divide the four now??

OpenStudy (anonymous):

or add d to 13??

OpenStudy (whpalmer4):

oh, don't be silly :-) \[9 = -d + 4d\]\[9 = 3d\]\[d=\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i got it...

OpenStudy (whpalmer4):

\[d = \] prove that you got it :-)

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

right. so now we know \(d = 3\), what is \(a_1=\)

OpenStudy (anonymous):

1,4,7,10,13

OpenStudy (anonymous):

an= 1+ (n-1)(3)

OpenStudy (whpalmer4):

right! so the process we used to solve that system of two equations in two unknowns is called substitution. we take one of the equations, solve it for one variable in terms of the other, and then substitute it into the second equation, giving us an equation in terms of just one variable. We solve that, and then use the first equation to find the value of the variable we substituted.